OpenStudy (anonymous):

*************Medals and fans************** What is the slope of the line whose equation is 2x – 4y = 10?

OpenStudy (anonymous):

Last one

OpenStudy (igreen):

Rearrange it into point-slope form: 2x - 4y = 10 Subtract 2x to both sides: -4y = -2x + 10 Divide -4 to both sides: \(y = \dfrac{1}{2}x + 2.5\) Now it's in the form of y = mx + b, where m = slope. So what's the slope gonna be? @The_Jokers_wife

OpenStudy (anonymous):

Slope of line in standard form ie \(ax + by = c\) is given by : \[Slope = \frac{-a}{b}\]

OpenStudy (anonymous):

ummm... 11

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

\(12\).

OpenStudy (igreen):

Lol^

OpenStudy (igreen):

y = mx + b m = slope y = 1/2x + 2.5 slope = ? @The_Jokers_wife

OpenStudy (anonymous):

Not funny @iGreen so 12 over 22

OpenStudy (anonymous):

He or she is saying that to me. @The_Jokers_wife

OpenStudy (anonymous):

Sorry, I did not see that igreen has asked you a question. No, it is not 11, not 12.

OpenStudy (anonymous):

Do you know how to compare two equations?

OpenStudy (anonymous):

Not really I really just forgot all this stuff in school and now I'm like what?????!!!!!

OpenStudy (anonymous):

Then why do you want to learn this stuff again? :P

OpenStudy (anonymous):

I'm in collage and I have to learn this stuff to pass my husband is going to pissed off

OpenStudy (anonymous):

See, I am comparing two equations, just try to feel the feelings I want to convey :P